OpenStudy (anonymous):

Triangles QVS and RTS are similar. Find side VS. QV is 8,RT is 4 and TS is 14

OpenStudy (anonymous):

Vs should equal 14 im not completely sure but if they are similar then Vs=Ts

OpenStudy (anonymous):

they are two triangles connected, then vs has a long arrow passing v,t, and s with a x on it

OpenStudy (anonymous):

how are they connected can you be more specific or can u send me a link

OpenStudy (anonymous):

its looks like that but with the numbers that i gave you

OpenStudy (campbell_st):

well you match the corresponding sides and write them as ratios this means \[\frac{VS}{TS} = \frac{VQ}{TR}\] using the values of the sides \[\frac{ x}{8} = \frac{10}{6}\] now you can solve for x

OpenStudy (anonymous):

no QV is 8,TS is 14, and RT is 4

OpenStudy (campbell_st):

the measurements you posted in the question don't match the measurements in the diagram

OpenStudy (anonymous):

obviously thats why i said that the picture is the same but not the numbers

OpenStudy (campbell_st):

well it's the same technique \[\frac{VS}{TS} = \frac{VQ}{RT}\] substitute your measurements \[\frac{x}{14} = \frac{8}{4}\] solve for x

OpenStudy (anonymous):

okay well i cross multiplied and got 4x=112, divided both by 4 and got x=28

OpenStudy (campbell_st):

that's what I got... I just looked at the right hand sides and simplified the fraction to 2/1 so larger measurement is double the smaller saves using algebra